The following table shows the State-of-Origin for the students in a large statistics class at Duke University.

State Number of Students
North Carolina 35
Virginia 9
South Carolina 8
Florida 15
Another country 8
Another state 24

(a) Determine the distribution for the State-of-Origin variable using proportions (relative frequencies) instead of counts (frequencies) as above. (Give each answer to 3 decimal places.)

State Proportion of Students
North Carolina
Virginia
South Carolina
Florida
Another country
Another state
Answer #1

total number of students (n) = 35+9+8+15+8+24 = 99

relative frequency = (individual count)/(total number n)

North carolina = (individual count)/(total number n) = 35/99 = 0.354

Virginia = (individual count)/(total number n) = 9/99 = 0.091

South carolina = (individual count)/(total number n) = 8/99 = 0.081

Florida = (individual count)/(total number n) = 15/99 = 0.152

Another country = (individual count)/(total number n) = 8/99 = 0.081

Another state = (individual count)/(total number n) = 24/99 = 0.242

State Proportion of Students
North Carolina 0.354
Virginia 0.091
South Carolina 0.081
Florida 0.152
Another country 0.081
Another state 0.242